Consider a project to supply 102 million postage stamps per year to the U.S. Postal Service for the next five years. You have an idle parcel of land available that cost $1,920,000 five years ago; if the land were sold today, it would net you $2,120,000 aftertax. The land can be sold for $2,320,000 after taxes in five years You will need to install $5.42 million in new manufacturing plant and equipment to actually produce the stamps; this plant and equipment will be depreciated straight-line to zero over the projects five-year life The equipment can be sold for $520,000 at the end of the project. You will also need $620,000 in initial net working capital for the project, and an additional investment of $52,000 in every year thereafter. Your production costs are 0.52 cents per stamp, and you have fixed costs of $1,070,000 per year. If your tax rate is 30 percent and your required return on this project is 10 percent, what bid price should you submit on the contract? (Do not round intermediate calculations and round your final answer to 5 decimal places. (e.g., 32.16161)) Bid price
Answer #1

As a first step let's calculate the series of cash flows for investment.

  1. Investment in land at t=0 will be opportunity cost of land today = - sale price of land today = - $ 2,120,000
  2. Return of investment in land at the end of year 5 = sale price of land at the end of year 5 = $ 2,320,000
  3. Investment in plant and equipment at t=0 will be -$ 5,420,000
  4. Salvage value at the end of year 5 = $ 520,000
  5. Tax basis at the end of year 5 = 0 (as asset will be fully depreciated by then)
  6. Gain on sale of plant & equipment = Salvage value - tax basis = 520,000 - 0 = 520,000
  7. Tax on gain = Tax rate, T x Gain on sale = 30% x 520,000 = 156,000
  8. Post tax salvage value = Salvage value - tax on gain on sale = 520,000 - 156,000 = 364,000
  9. Initial investment in working capital = -620,000
  10. Subsequent investment in working capital per year for next four year = - 52,000
  11. Release of working capital at the end of year 5 = cumulative investment in working capital = 620,000 + 4 x 52,000 = $ 828,000

Discount rate, R = 10%

PV factor for year N = (1 + R)(-N)

Let's put all these into a table. Please see the table below:

Tax rate

T

30%

Year, N

0

1

2

3

4

5

Investment in land

A

(2,120,000)

2,320,000

Investment in new plant & equipment

B

(5,420,000)

Post tax Salvage value of plant & equipment

C

364,000

Investment in working capital

D

(620,000)

(52,000)

(52,000)

(52,000)

(52,000)

828,000

Net cash flows towards investment

E = A + B + C + D

(8,160,000)

(52,000)

(52,000)

(52,000)

(52,000)

3,512,000

Required return

R

10%

PV factor

(1+R)^(-N)

1.0000

0.9091

0.8264

0.7513

0.6830

0.6209

PV of net investment

F = E x PV factor

(8,160,000)

(47,273)

(42,975)

(39,068)

(35,517)

2,180,676

NPV of investments

Sum of all F

(6,144,157)

Sum of PV factors from year 1 to 5

Sum of PV factor for years 1 to 5 only

3.7908

NPV of all the investments = -6,144,157

Sum of PV factors for year 1 to 5 = 3.7908

Now, let's say S is the bid price per stamp in $. Then annual operating post tax cash flow will be:

[(Sale Price - production cost ) x Nos. of stamps to be supplied annually - Fixed Cost - Depreciation] x (1 - Tax rate) + Depreciation

Production cost = 52 cents = $ 0.0052

Nos. of stamps to be supplied annually = 102 mn = 102,000,000

Fixed cost per annum = 1,070,000

Annual depreciation = Cost of Plant & machinery / Life = 5,420,000 / 5 = 1,084,000

Annual Post tax operating cash flow = [(S - 0.0052) x 102,000,000 - 1,070,000 - 1,084,000] x (1 - 30%) + 1,084,000 = 71,400,000 x S - 795,080

Sum of PV of all annual operating cash flows over year 1 to 5 = Annual Operating cash flow x Sum of PV factor over 5 years = (71,400,000 x S - 795,080) x 3.7908

If we bid at a price such that NPV of the project is 0 then

Sum of PV of all annual operating cash flows over year 1 to 5 + NPV of investments = 0

Hence, (71,400,000 x S - 795,080) x 3.7908 - 6,144,157 = 0

Hence, S = $ 0.03383604

So, please enter a value of 0.03384 (rounded off to 5 places of decimal)